OpenStudy (anonymous):

Graph the linear equation. y = 3x – 4

OpenStudy (anonymous):

Assume that a Value of x And put this in the equation above to get the value of y now plot the point (x,y) on the graph .... In example, Lets assume x equals 1 then y = -3x -4 => y = -3(1) - 4 = -3-4 = -7 so plot point (1,-7) also as this is a linear equation it represents a st line And any st. line can be defined by only two points So, now assume x equals 2 or -1 (any value you want really but lets say -1) So x=-1 put in the equation to get y y = -3x -4 = -3*(-1) -4 = 3-4 = -1 so point is (-1,-1) Now plot the points P1(1,-7) and P2(-1,-1) And join the two points (P1 & P2) and you get your graph for this problem. Hope I helped.